OpenStudy (anonymous):

Find all values of x; tan x=1/√3

OpenStudy (anonymous):

see when tan x = 1/sqrt(3) tanx = tan (pi/6) now when tanx = tant then x= n*(pi)+t where n is any integer(0,+-1,+-2 ,+-3,....)

OpenStudy (anonymous):

here t= pi/6

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

in particular u can substitute values for n

OpenStudy (anonymous):

Oh btw, how did you get pi/6